A 100 nC charge is placed at the origin. A -50 nC charge is placed at x= 5 cm and a 200 nC is placed at x= -10 cm.

 a. What is the net electrostatic force acting on 100 nC charge?


1
Expert's answer
2022-03-11T09:03:41-0500

"F_{net}=F_1+F_2"

"F=\frac{kq_1q_2}{r^2}+\frac{kq_2q_3}{r^2}"

"F=kq_2[\frac{200\times10^{-9}}{100\times10^{-4}}-\frac{50\times10^{-9}}{25\times10^{-4}}]"


"F=9\times10^9\times100\times10^{-9}[\frac{200\times10^{-9}}{100\times10^{-4}}-\frac{50\times10^{-9}}{25\times10^{-4}}]=0.04N"